OpenStudy (anonymous):

Please Help. Im just having trouble remembering how to do this. Calculate the average rate of change of the given function over the given interval f(x) = x^2 − 3;[8, 10] Answer is 18

OpenStudy (anonymous):

\[\frac{f(10)-f(8)}{10-8}\] is what you are looking for

OpenStudy (anonymous):

since the 3's will add up to zero, it is really a matter of computing \[\frac{10^2-8^2}{2}=\frac{36}{2}=18\]

OpenStudy (anonymous):

ok thank you!